OpenStudy (anonymous):

REALLY NEED SOME HELP!!! (all of these questions are from a rational numbers quiz, please help!) 1.Simplify: Negative four and one-fifth minus negative thirteen and one-tenth A) Eight and Nine Tenths B) Negative eight and Nine tenths C) Negative Seventeen and Three tenths D) Seventeen and Three Tenths 2.Use absolute value to express the distance between −12 and −15 on the number line. A) |-12 - (-15)| = -27 B) |-12 - (-15)| = -3 C) |-12 - (-15)| = 3 D) |-12 - (-15)| = 27 Third question will be in the comments!!!!

OpenStudy (anonymous):

Arlene is showing her work in simplifying −7.9 + 8.2 − 3.4 + 2.1. Identify any errors in her work or in her reasoning. Write feedback to Arlene explaining the mistake she made and how to correct her work.

OpenStudy (johnweldon1993):

\[\large -4\frac{1}{5} - (-13\frac{1}{10})\] \[\large -4\frac{1}{5} + 13\frac{1}{10}\] Change into improper fractions \[\large \frac{-21}{5} + \frac{131}{10}\] Common denominator \[\large \frac{-42}{10} + \frac{131}{10}\] \[\large \frac{-42 + 131}{10}\] \[\large \frac{89}{10} = 8\frac{9}{10}\]

OpenStudy (johnweldon1993):

If you have any questions just ask :)

OpenStudy (anonymous):

Thank you!!! Do you know the answer to the other two questions??

OpenStudy (johnweldon1993):

Oh sorry hun lol looked away for a second...gimme 1 sec :)

OpenStudy (johnweldon1993):

So absolute value tells you how far apart numbers are *always positive since it is just a distance ex: From 1 to 4 = |4 - 1| = |3| = 3 From -1 to -4 = |-4 - (-1)| = |-4 + 1| = |-3| = 3 <--since the absolute value turns everything positive at the end...so it looks like choice C here for you :)

OpenStudy (anonymous):

Yay!! Thank you!!!

OpenStudy (anonymous):

Oops, I forgot about the question in the comments..do you know how to do that one?

OpenStudy (johnweldon1993):

Haha I just saw that too Okay 1 sec

OpenStudy (johnweldon1993):

There is no mistake...Arlene is good XD

OpenStudy (anonymous):

Thank you very much!! You're awesome!

OpenStudy (johnweldon1993):

No problem :)